# The Exchangeability Assumption for Permutation Tests of Multiple Regression Models

### v1

@@ Abstract
Permutation tests are a powerful and flexible approach to inference via resampling. As computational methods become more ubiquitous in the statistics curriculum, use of permutation tests has become more tractable. At the heart of the permutation approach is the exchangeability assumption, which determines the appropriate null sampling distribution. We explore the exchangeability assumption in the context of permutation tests for multiple linear regression models. Various permutation schemes for the multiple linear regression setting have been previously proposed and assessed in the literature. As has been demonstrated previously, in most settings, the choice of how to permute a multiple linear regression model does not materially change inferential conclusions. Regardless, we believe that (1) understanding exchangeability in the multiple linear regression setting and also (2) how it relates to the null hypothesis of interest is valuable. We also briefly explore model settings beyond multiple linear regression (e.g., settings where clustering or hierarchical relationships exist) as a motivation for the benefit and flexibility of permutation tests. We close with pedagogical recommendations for instructors who want to bring multiple linear regression permutation inference into their classroom as a way to deepen student understanding of resampling-based inference.
@@ 1 Introduction
Statistical inference is based on modeling the variability inherent in a dataset. Many of the models and analysis methods taught in a standard undergraduate statistics curriculum rely on asymptotic normal theory, with the theoretical underpinnings driven by the Central Limit Theorem. However, permutation tests are becoming increasingly popular because they provide a flexible approach for a wide scope of problems. Unlike methods base on the Central Limit Theorem, permutation tests do not generally require distributional or sample size assumptions. They do, however, require exchangeability, an idea which comprises much of the substance of the paper and which will be introduced in Section 1.1.
Permutation tests were among the first inferential tests conceived and used widely (Fisher 1935), in the context of categorical data. Permutation tests have since been expanded to cover many different modeling contexts and are often presented as a way to deepen an understanding of sampling distributions and normal theory methods (Ismay and Kim 2020; Chance and Rossman 2021; Baumer, Kaplan, and Horton 2024; Çentinkaya-Rundel and Hardin 2023). Normal inference with simple least squares regression requires that the data are independent, follow a linear model, and have error terms which are normal with equal variance. Permutation tests allow for inference on the least squares model where the normality condition can be dropped. The computational application of a permutation test in the simple linear model case requires the analyst to permute, equivalently, either the explanatory or the response variable before re-fitting the least squares model under the null hypothesis. However, the simple linear regression model is somewhat limited in the types of problems it can handle, and there has been a recent push to infuse the statistics curriculum with multivariate thinking (Carver et al. 2016).
@@
Myriad approaches to permutation approaches in the multiple linear regression (MLR) model have been proposed in the literature. Manly (1997) and Winkler et al. (2014) summarize and synthesize a variety of MLR permutation methods. In N. Draper and Stoneman (1966), the treatment variable of interest is shuffled, while Manly (1986) shuffles the response variable itself. Permuting model residuals are considered in multiple papers. Freedman and Lane (1983) permute residuals from a null model, ter Braak (1992) permutes residuals from a full model, Kennedy (1995) permutes values from a model that residualizes both the response variable and the treatment variable, and Huh and Jhun (2001) permutes under conditions of a block structure experimental design. Still and White (1981) consider a special ANOVA case with interaction.
MLR provides a foundation for multivariate thinking that is both accessible to undergraduates and often taught in many introductory statistics classes and second courses in applied statistics. Permutation methods are ubiquitous in graduate studies, where student build up intuition, theory, and computation. Some aspects of permutation testing in the MLR setting are not fully straightforward. As an example, the “permute” step of a permutation test in the MLR is not as obvious as in the simple linear regression case, with the determination of how to permute depending on the underlying structure of the data, the specific hypotheses begin tested, and the assumption of exchangeability.
@@
While not necessarily straightforward, working through some of the permutation options for MLR affords an understanding of exhangeability and permutation tests more generally that sets up students for understanding permutation tests in more complex settings. We review some common MLR permutation methods and motivate how and why exploration of permutation-based inference for multiple regression models is valuable.
@@ 1.1 Exchangeability
At the foundation of valid permutation tests is the condition of exchangeability (Pitman 1937; Good 2002). While there are typically no distributional or sample size restrictions on a permutation test, how to incorporate the exchangeability condition can sometimes be difficult (Welch 1990). D. Draper et al. (1993) gives a detailed account of the issues of exchangeability, including when more complicated settings where the data have serial correlation.
@@
Data are exchangeable under the null hypothesis if the joint distribution from which the data came is the same before permutation as after permutation when the null hypothesis is true.
@@
Note, for example, if a confounding effect is present in a MLR model, the data are often not exchangeable, even when the null hypothesis is true. Consider an example whose goal is to model performance in college as the response variable ($Y$) based on whether the student took Advanced Placement (AP) courses in high school ($X_{2}$, the treatment variable) and the student’s family income level ($X_{1}$). It has been well documented that work on AP exams is associated with socio-economic status (Kolluri 2018). Assume that the null hypothesis is true; that is, having taken the AP exam is not linearly related to performance in college after controlling for family income.

### v2

@@ Abstract
Permutation tests are a powerful and flexible approach to inference via resampling. As computational methods become more ubiquitous in the statistics curriculum, use of permutation tests has become more tractable. At the heart of the permutation approach is the exchangeability assumption, which determines the appropriate null sampling distribution. We explore the exchangeability assumption in the context of permutation tests for multiple linear regression models, including settings where the assumption is not tenable. Various permutation schemes for the multiple linear regression setting have been proposed and assessed in the literature. As has been demonstrated previously, in most settings, the choice of how to permute a multiple linear regression model does not materially change inferential conclusions with respect to Type I errors. However, some violations (e.g., when clustering is not appropriately accounted for) lead to issues with Type I error rates. Regardless, we believe that understanding (1) exchangeability in the multiple linear regression setting and also (2) how it relates to the null hypothesis of interest is valuable. We close with pedagogical recommendations for instructors who want to bring multiple linear regression permutation inference into their classroom as a way to deepen student understanding of resampling-based inference.
@@ 1 Introduction
Statistical inference is based on modeling the variability inherent in a dataset. Many of the models and analysis methods taught in a standard undergraduate statistics curriculum rely on asymptotic normal theory, with the theoretical underpinnings driven by the Central Limit Theorem. However, permutation tests are becoming increasingly popular because they provide a flexible approach for a wide scope of problems. Unlike methods based on the Central Limit Theorem, permutation tests do not generally require distributional or sample size assumptions. They do, however, require exchangeability, an idea which comprises much of the substance of our paper and which will be introduced in Section 1.1.
Permutation tests were among the first inferential tests conceived and used widely (Fisher 1935) in the context of categorical data. Permutation tests have since been expanded to cover many different modeling contexts and are often presented as a way to deepen an understanding of sampling distributions and normal theory methods. Permutation methods have long been pervasive in the statistics curriculum for graduate studies, where students build up intuition, theory, and computation. However, the undergraduate statistics curriculum has historically not had space for extended explication of permutation methods, although many modern introductory textbooks do introduce the basics of permutation tests (Ismay and Kim 2020; Chance and Rossman 2021; Baumer, Kaplan, and Horton 2024; Çentinkaya-Rundel and Hardin 2023). Regression, on the other hand, plays a central role within undergraduate minor and major programs (American Statistical Association 2014).
Normal inference with simple least squares regression, also known as simple linear regression (SLR), requires that the data are independent, follow a linear model, and have error terms which are approximately normal with equal variance. Permutation tests are attractive, since they allow for inference on the least squares model where the normality condition does not hold (particularly with small sample sizes, since it may not be tenable to invoke the Central Limit Theorem).
The computational application of a permutation test in the SLR model case requires the analyst to permute, equivalently, either the predictor or the outcome variable before re-fitting the least squares model. (A third option, permuting the residuals, is discussed later in Section 3.3.3 and Section 3.3.4.)
However, the SLR model is somewhat limited in the types of problems it can handle, and there has been a recent push to infuse the statistics curriculum with multivariate thinking (Carver et al. 2016). Multiple linear regression (MLR) provides a foundation for multivariate thinking that is both accessible to undergraduates and often taught in many introductory statistics classes and second courses in applied statistics (see, for example, a discussion of the undergraduate statistics curriculum in which teaching exchangeability is explicitly referenced in Kennedy-Shaffer (2024)).
Unlike SLR, in the MLR model, what to permute is not immediately obvious within the inferential process. The “permute” step of a permutation test in the MLR model is not as obvious as in the SLR case, with the determination of how to permute depending on the underlying structure of the data, the specific hypotheses begin tested, and the assumption of exchangeability.
@@
The statistics literature includes myriad proposals for different permutation approaches in the MLR model. Manly (1997) and Winkler et al. (2014) summarize and synthesize a variety of MLR permutation methods. In N. Draper and Stoneman (1966), the treatment variable of interest is shuffled, while Manly (1986) shuffles the outcome variable itself. Multiple authors consider permuting model residuals: Freedman and Lane (1983) permute residuals from a null model, ter Braak (1992) permutes residuals from a full model, Kennedy (1995) permutes values from a model that residualizes both the outcome variable and the treatment variable, and Huh and Jhun (2001) permute under conditions of a block structure experimental design. Still and White (1981) consider a special ANOVA case with interaction. Recent work has described a variety of robust permutation approaches (DiCiccio and Romano 2017; Helwig 2019a, 2019b).
@@
While not necessarily straightforward, working through some of the permutation options for MLR provides an understanding of exchangeability, and permutation tests more generally, that sets up students for understanding permutation tests in more complex settings. Our goal is to present MLR permutation methods, explore additional complications, and motivate how and why an investigation into permutation-based inference for multiple regression models is valuable as a way to expand understanding of inference and statistical foundations.
@@ 1.1 Exchangeability
At the foundation of valid permutation tests is the condition of exchangeability (Pitman 1937; Good 2002). While there are typically no distributional or sample size restrictions on a permutation test, recognizing how to incorporate the exchangeability condition can sometimes be difficult (Welch 1990). D. Draper et al. (1993) gives a detailed account of the issues of exchangeability, including more complicated settings, such as serial correlation.
@@
Data are exchangeable under the null hypothesis if the joint distribution from which the data came is the same before a permutation as after a permutation when the null hypothesis is true.
For linear regression models with a single predictor (SLR), equivalent results will ensue no matter whether the outcome variable ($Y$) or the predictor variable ($X$) is permuted.
Carrying out a permutation procedure involves calculation of a statistic for each of the permuted samples, then comparing the observed statistic to the permuted distribution where the association between $Y$ and $X$ has been broken. Here it doesn’t matter whether the outcome $Y$ or the predictor $X$ is permuted: the results will be equivalent.
A typical choice of statistic is a t-statistic, where the estimated regression coefficient is divided by the estimated standard error (Janssen 1997; Konietschke and Pauly 2012). Such a pivotal statistic has attractive properties, including providing robustness to modest deviations from exchangeability.
Undertaking a permutation test is more complicated when a second predictor is of interest, thereby making the MLR the desired model.
1.2 Motivating example
@@
Consider an example whose goal is to model performance in college as the outcome variable ($Y$) based on whether the student took Advanced Placement (AP) courses in high school ($X_{1}$, the “treatment” variable) and the student’s family income level ($X_{2}$, a potential confounder). It has been well-documented that work in AP courses is positively associated with socio-economic status (Kolluri 2018). Assume that the null hypothesis ($\beta_{1}=0$) is true; that is, having taken an AP course is not linearly related to performance in college after controlling for family income. For the moment, we assume that the families are independent and only one student is included from each family.
Understanding exchangeability is sometimes best understood when it is violated. Imagine now carrying out a test from the given example setting by permuting $Y$ for this MLR model. The permutation process would break the relationship between $X_{1}$ and $Y$ (as desired), would preserve the relationship between $X_{1}$ and $X_{2}$ (as desired), but would also unfortunately break the relationship between $Y$ and $X_{2}$.